OpenStudy (anonymous):

Which are the solutions to the given equation? Answers are rounded to the nearest hundredth when necessary. x^2-2x-1=0

OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

add one get \[x^2-2x=1\]complete the square get \[(x-1)^2=1+1=2\] so \[(x-1)^2=2\\ x-1=\pm\sqrt2\\ x=1\pm\sqrt2\]

OpenStudy (haseeb96):

just use the factorization method and have a look on @misty1212 explanation